OpenStudy (anonymous):

For what values of x does the series \(\sum_{n=1}^\infty \frac{(4x-5)^{2n+1}}{n^{\frac{3}{2}}}\) converge absolutely? Conditionally?

OpenStudy (anonymous):

is the denominator \[\sqrt{n^3}\]?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

let me try again

OpenStudy (anonymous):

For the interval of convergence I got 1 < x < 3/2, if that helps. I'm also not sure how to analyze the convergence/divergence at the endpoints.

OpenStudy (anonymous):

i am being foolish you need ratio test for this one, should get the limit is 1 so you have to solve \[|4x-5|<1\]

OpenStudy (anonymous):

oh ok you are right

OpenStudy (anonymous):

so lets replace x by 1 and see what happens

OpenStudy (anonymous):

you get \[\sum\frac{(-1)^{2n+1}}{\sqrt{n^3}}\]which certainly converges because it is alternating and the terms go to zero it also converges absolutely because the exponent on the denominator is greater than one

OpenStudy (anonymous):

replace x by 3/2 and you get \[\sum\frac{1}{\sqrt{n^3}}\] which also converges for the same reason above so it converges at the endpoints of the interval as well

OpenStudy (anonymous):

How do you determine what values of x converge conditionally?

OpenStudy (anonymous):

ok first off if it converges absolutely it converges conditionally

OpenStudy (anonymous):

you already have the interval of convergence, you computed it as \[1<x<1.5\] now at the left hand endpoint you get an alternating sequence, so it is possible that it converges "conditionally" meaning that the absolute value does not converge. but in this case the absolute value does converge because the exponent in the denominator is greater than one